A pair of narrow, parallel slits separated by 0.245 mm is illuminated by the green component from a mercury vapor lamp (λ = 546.1 nm). The interference pattern is observed on a screen 1.23 m from the plane of the parallel slits.

(a) Calculate the distance from the central maximum to the first bright region on either side of the central maximum.
 mm

(b) Calculate the distance between the first and second dark bands in the interference pattern.
 mm
Answer #1

a)
The first maxima after the central maxima occurs at d sin\theta = \lambda
Where d is the separation between the slits and \lambda is the wavelength.
sin\theta = \lambda/d
= (0.546 x 10-6) / (0.245 x 10-3)
= 2.23 x 10-3.
Since \theta is very small, sin\theta = tan\theta
tan\theta = y/D
Where y is the distance from the central maxima to the first maxima and D is the distance to the screen.
y = D tan\theta
= 1.23 x 2.23 x 10-3
= 2.74 x 10-3 m
= 2.74 mm.

b)
For the first dark band, d sin\theta1 = \lambda/2
sin\theta1 = 0.5\lambda / d
For the second dark band, d sin\theta2 = 3\lambda/2
sin\theta2 = 1.5\lambda / d

Consider y2 and y1 be be distance to the second and first dark band respectively.
y2 - y1 = D [tan\theta2 - tan\theta1]
= D [sin\theta2 - sin\theta1]
= D/d[1.5\lambda - 0.5 \lambda]
= D\lambda/d
= [1.23 x 546.1 x 10-9] / [0.245 x 10-3]
= 2.74 x 10-3 m.
= 2.74 mm

Note: The distance between two successive bright fringe of dark fringe is the same.